the last digit of 2137^754 is.....


a) 1

b)3

c)7

d) 9

Grade:12

5 Answers

sahil malhotra
13 years ago

8

mihir mehan
13 years ago

bhaiya yeh toh option hi nahi hai !!!!!!!!!  please also explain how you reached that answer whenever you post a reply.

Askiitians Expert Sahil Arora - IITD
13 years ago
Dear mihir
2137^754=(2130+7)^754
on binomial expansion,we find that the first teram is not divisible by 10 while others are divisible by 10..
SO only first term will contribute to last digit..
the first term is 7^754
the last digit of this will be the last digit of given question.'
now observe the last digit in following arguments.
7^1=7,
7^2=49,
7^3=513
7^4=last digit being 1
and then the last digits repeat
the last digits are 7,9,3,1,7,9,3,1,
SO 7^(4n+1)---last digit is 7
7^(4n+2)---- last digit is 9
so on dividing 754 by 4 we get 2 as remainder...
so the last digit will be 9...
Please feel free to post as many doubts on our discussion forum as you can. If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very quickly. We
are all IITians and here to help you in your IIT JEE preparation.
All the best mihir !!!
Regards,
Askiitians Experts
Sahil Arora
IIT Delhi
RAJESH
9 years ago
units of 7754=7377*72​​                                                                                                                          therefore 72=49 and the units digit of it is 9
Aakash Yadav
7 years ago
Since we can see the 7^1=7 7^2=49 7^3=343=7^4has it`s unit digit 1 so now see for 752 the unit digit=1 now for 753unitdigit=7 754unit digit =9